OpenStudy (anonymous):

The function f:x = 4-3sinx is defined for the domain 0

OpenStudy (anonymous):

sinx varies from -1 to 1

OpenStudy (anonymous):

so, 4-3sin x varies from 1 to 7 so, any real value of k except 1 to 7 has to be eliminated so, \[k \in R-[1,7]\]

OpenStudy (anonymous):

sorry, any real value between 1 to 7 has to be eliminated^^

OpenStudy (anonymous):

It's confusing. Can you explain?

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

Is the answer -1 to 1 or 1 to 7?

OpenStudy (anonymous):

k has any real value except 1 to 7

OpenStudy (anonymous):

But it has values for 1-7

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

I'll close this and open another, but please still try and help!

OpenStudy (anonymous):

the lowest value of the function= 4-3*1=1 highest value of the function= 4-3*(-1)=7 so, if k is between 1 to 7, the equation will have solution but it is told for which values of k it will have NO SOLUTION so, any value except 1 to 7

OpenStudy (anonymous):

got it? @order ?

OpenStudy (anonymous):

Yes, I got it.